# Toward Foundation Models for Online Complex Event Detection in CPS-IoT: A Case Study

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Machine learningccs: Computer systems organization Embedded and cyber-physical systemsccs: Artificial intelligence Knowledge representation and reasoningccs: Artificial intelligence Hybrid symbolic-numeric methods
@@ 6.3. Evaluation
We evaluate model performance across different training set sizes, as shown in Fig. 3. The results indicate that Mamba achieves the best performance, followed by LSTM. The Neural + X models underperform compared to end-to-end models, likely due to errors and noise introduced by the Neural AE classifier. This also explains why Neural AE + FSM performs worse despite incorporating correct human-defined CE rules. Additionally, we test model generalization on out-of-distribution (OOD) complex events lasting 15 and 30 minutes, which follow the same CE rules but extend their temporal spans. As shown in Fig. 4, Mamba exhibits the best generalization among both end-to-end neural models and Neural AE + FSM. Table 3 further shows that Mamba, when trained with more labeled sensor data on 5-minute sequences, improves generalization to longer unseen traces, suggesting that it efficiently learns CE rules as training data increases. These findings indicate that Mamba may serve as a strong backbone for foundation models targeting long and complex temporal reasoning.
Figure 3. Positive $F1$ scores of models on complex events with different training data.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: The 2nd International Workshop on Foundation Models for Cyber-Physical Systems & Internet of Things ; May 6–9, 2025; Irvine, CA, USAbooktitle: The 2nd International Workshop on Foundation Models for Cyber-Physical Systems & Internet of Things (FMSys ’25), May 6–9, 2025, Irvine, CA, USAdoi: 10.1145/3722565.3727198isbn: /2025/05ccs: Computing methodologies Machine learningccs: Computer systems organization Embedded and cyber-physical systemsccs: Artificial intelligence Knowledge representation and reasoningccs: Artificial intelligence Hybrid symbolic-numeric methods
@@ 6.3. Evaluation
We evaluate model performance on CEs defined in Table 1 across different training set sizes, as shown in Fig. 3. The results indicate that Mamba achieves the best performance, followed by LSTM. The Neural + X models underperform compared to end-to-end models, likely due to errors and noise introduced by the Neural AE classifier. This also explains why Neural AE + FSM performs worse despite incorporating correct human-defined CE rules. Additionally, we test model generalization on out-of-distribution (OOD) complex events lasting 15 and 30 minutes, which follow the same CE rules but extend their temporal spans. As shown in Fig. 4, Mamba exhibits the best generalization among both end-to-end neural models and Neural AE + FSM. Table 3 further shows that Mamba, when trained with more labeled sensor data on 5-minute sequences, improves generalization to longer unseen traces, suggesting that it efficiently learns CE rules as training data increases. These findings indicate that Mamba may serve as a strong backbone for foundation models targeting long and complex temporal reasoning.
Figure 3. Boxplot of positive $F_{1}$ scores on complex events with different training data sizes. Solid line in the box shows median; dashed line in the box shows mean.